At a certain store, the price of a calculator is $150, the price of a radio is $75, and the price of a printer is 16 percent less than the price of a radio. By what percent is the price of a printer less than the price of a calculator?
a) 39%
b) 58%
c) 63%
d) 87%

Answer by Theo(13342)

price of a calculator is 150.
price of a radio is 75.
price of a printer is 16% less than the price of a radio.

the price of the printer is 75 minus .16 * 75 = .84 * 75 = 63.

the price of the calculator minus the price of the printer is 150 minus 63 = 87.
87 / 150 = .58 = 58%.
the price of the printer is 58% less than the price of the calculator.
the price of the printer is 150 minus .58 * 150 = 63.
this confirms the solution is correct.

the solution is 58% which is selection b.

let me know if you have any questions.
theo